Consider a 15-year, $140,000 mortgage with a rate of 5.80 percent. Five years into the mortgage, rates have fallen to 5 percent. What would be the monthly saving to a homeowner from refinancing the outstanding mortgage balance at the lower rate
Answer #1
Step-1:Monthly payment at the beginning of mortgage period
Monthly payment =pmt(rate,nper,pv) Where,
$       1,166.33 rate = 5.80%/12 = 0.004833333
nper = 15*12 = 180
pv = $ -1,40,000.00
Step-2:Mortgage balance after 5 years
Mortgage balance =-pv(rate,nper,pmt) Where,
$ 1,06,011.49 rate = 5.80%/12 = 0.004833333
nper = (15-5)*12 = 120
pmt = $         1,166.33
Step-3:Revised monthly payment
Monthly payment =pmt(rate,nper,pv) Where,
$       1,124.42 rate = 5.00%/12 = 0.004166667
nper = (15-5)*12 = 120
pv = $ -1,06,011.49
Step-4:Saving of monthly payment
Saving of monthly payment = Monthly payment at beginning - Monthly payment after 5 years
= $       1,166.33 - $ 1,124.42
= $             41.91